OpenStudy (anonymous):

Which statement about the following equation is true? 3x2 - 8x + 5 = 5x2

OpenStudy (anonymous):

The discriminant is less than 0, so there are two real roots. The discriminant is greater than 0, so there are two real roots. The discriminant is less than 0, so there are two complex roots. The discriminant is greater than 0, so there are two complex roots.

OpenStudy (anonymous):

3x^2 - 8xt5=5x^2

OpenStudy (anonymous):

please help!!!!!!

OpenStudy (anonymous):

pleaase!!!!!!!!!!!!!

OpenStudy (kamibug):

Here's what I know about discriminants from my notes. I hope it helps. :) Formula to find the discriminant: b^2 - 4ac If the discriminant is 0 there will be 1 rational solution. If the discriminant is a positive perfect square there will be 2 rational solutions. If the discriminant is a positive integer, not a perfect square, there will be 2 irrational solutions. If the discriminant is a negative, there will be 2 complex solutions.

OpenStudy (kamibug):

Can you show me the equation you were given from the question using the equation tool provided under the chat? :)

OpenStudy (anonymous):

3x2 – 8x + 5 = 5x2

OpenStudy (anonymous):

i tried but nothing popped up from it

OpenStudy (anonymous):

its a - where the question marks are

OpenStudy (anonymous):

what do you think it is

OpenStudy (kamibug):

It's alright. :) Well we want the equation to look like this ... ax^2 + bx + c = 0 ... So what we gotta do now is set it equal to zero by subtracting 5x^2 from both sides. 5x^2 and 3x^2 are like terms. So when we simplify this will be our new equation ... -2x^2 - 8x + 5 = 0 ... Now that we have it set like this we want to identify the numbers we will need for the discriminant formula. :) a: -2 b: -8 c: 5 Just plug those numbers into the formula. :) Got it?

OpenStudy (anonymous):

yeah ty